OpenStudy (anonymous):

28 men can finish the job in 60 days. At the start of the 16 day, 5 men were laid off and after the 45th, 10 more men were hired. How many days were delayed in finishing the job? Please explain, step by step.

OpenStudy (anonymous):

really? i am going to change my name to i hate math

OpenStudy (anonymous):

i am going to suggest this although there is probably a snappier way. 28 men in 60 days to one job so each man and do \[\frac{1}{28\times 60}=\frac{1}{1680}\] of the job per day

OpenStudy (anonymous):

then the equation becomes \[\frac{15\times 28}{1680}+\frac{23\times 30}{1680}+\frac{33\times x}{1680}=1\]

OpenStudy (anonymous):

or if you prefer \[33x+1110=1680\] and thus \[x=\frac{190}{11}=17.27\]so the total time was 45+17.27=62.27 and therefore the delay was 2 + (say 3) days

OpenStudy (anonymous):

from where did you base 33x+1110=1680?